OpenStudy (anonymous):

Find the slope-intercept form of the equation of the line passing through the points. (5/8 , 4/3) , (-7/2 , 1/6)

OpenStudy (anonymous):

can you start with the slope y2 - y1 ------- x2 - x1

OpenStudy (anonymous):

i got -7/6 / 23/8

OpenStudy (karatechopper):

ok then simplfy that^

OpenStudy (anonymous):

we can simplify this to -7/6 times 8/23 = -28/69

OpenStudy (anonymous):

Now we can use the point-slope formula y - y1 = m(x - x1)

OpenStudy (karatechopper):

pick any point and plug it in with your slope